please help me to solve this problem-
A machine makes electric bulb,in which 3 out of every 65 are found to be faulty.it made 45 faulty bulbs in a day.how many good bulbs were made that day?

Answer by dkppathak(439)

A machine makes electric bulb,in which 3 out of every 65 are found to be faulty.it made 45 faulty bulbs in a day.how many good bulbs were made that day?
3 out of 65
45 faulty means 45/3= 15 lot of 65 bulb
total bulb made by machine on that day =15x65=975 these include 45 faulty also
good bulb =975-45=930
or
(65-2)x15=63x15= 930 bulb
answer 930 bulb